It states "Amazon offers an interesting deal: (234 diapers for $46.99/ 160 diapers $34.99) if you pay $79 per year to join their "Amazon Mom" club, you get 20% off all diaper purchases.
What would the $46.99 box of diapers cost with that discount? Write an equation that provides the total cost C of buying x boxes of diapers in one year with the Amazon Mom membership. Use the equation to find the cost of buying 10 boxes of diapers in a year if you join Amazon Mom.

Answer by Boreal(15235)

46.99 -20% discount is 46.99*0.8=$37.59
C=79+(37.59)10=$79+$375.90=$454.90
or, for the other
C=79+27.99(10)=$79+279.90=$358.90
It isn't clear which 10 boxes, and it could even be some of each.
That would be C=79+37.59(x)+27.99(10-x), where x is the larger diaper box.